OpenStudy (anonymous):

find the remainder when 4^44 is divided by 44?

OpenStudy (anonymous):

2^88/44=?

OpenStudy (anonymous):

\[\frac{ 2^{88} }{ 44 } = \frac{ 2^2}{4} \frac{2^{86} }{ 11 }\]